# Cahn-Hilliard model with Schlögl Reactions: interplay of equilibrium and non-equilibrium phase transitions. II. Memory effects

### v1

@@ III The parametric dependence of solution
Naturally, for $\tau=0$ the equations (2.15), (2.17) coincide with corresponding equations of Part I [1] too. We denote the values, corresponding to $\tau=0$ as $\,\delta_{0}\,\,$ and $s_{0}$:
$\delta_{0}=\,\,\frac{3}{2}u_{2}-\left[\left(\frac{u_{2}}{2}-u_{1}\right)-\kappa\frac{\alpha u_{2}}{2}\right]\eta;$ (3.4)
@@
$\delta=\delta_{0}+\tau\delta_{1}\,\,\,\,s=s_{0}+\tau s_{1}.$ (3.6)
@@
Naturally, the corrections due to memory are proportional to the velocity of the front; there should be no corrections for static front.
The constraints (2.15) and (2.17) impose some limitations on the parameters $\delta$ and $s$ of the equilibrium phase transition, see (1.2), i.e. on the roots $\tilde{u}_{1},\,\,\tilde{u}_{2},\,\,\tilde{u}_{3}$ of the cubic equation
@@
These roots are positions of the extrema of the homogeneous part of the thermodynamic potential $\Phi\left(u\right)$; we denote the roots, corresponding to stable minima $\,\tilde{u}_{2},\,\,\tilde{u}_{3}$, and $\tilde{u}_{1}$ – corresponding to unstable maximum. The root $\tilde{u}_{3}=0$ coincides with one of the stationary states of the reaction system. The expressions for two remaining roots $\tilde{u}_{1}$ and $\,\tilde{u}_{2}$ yield the constraints imposed on the stationary values for the equilibrium transformation:
@@
We are mainly interested how the actual final concentration $u_{2}$ after the transition deviates from the equilibrium value $\,\tilde{u}_{2}$. Usually $\tau$ is presumed to be a small parameter; expanding (3.10) and keeping only linear in $\tau$ terms, we have
@@
Then the deviation of the equilibrium value from actual after transition is
$\tilde{u}_{2}-u_{2}=\frac{1}{2}\left\{\left[\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right]+\left(\delta_{1}+\frac{\delta_{0}\delta_{1}+2s_{1}}{\sqrt{\delta_{0}^{2}+4s_{0}}}\right)\tau\right\}.$ (3.12)
The $\tau$-independent term was calculated and discussed in [1]; here we are interested in the relative importance of memory effect
${\tilde{u}_{2}-u_{2}=\frac{1}{2}\left[\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right]}$
×${\times\left\{1+\left[\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right]^{-1}\times\left(\delta_{1}+\frac{\delta_{0}\delta_{1}+2s_{1}}{\sqrt{\delta_{0}^{2}+4s_{0}}}\right)\tau\right\}}.$ (3.13)
So we define
$F=\left[\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right]^{-1}\left(\delta_{1}+\frac{\delta_{0}\delta_{1}+2s_{1}}{\sqrt{\delta_{0}^{2}+4s_{0}}}\right).$ (3.14)
We will study the behavior of $F$ as the function of the system parameters. Following [1], we consider several special cases. First we consider CHS, i.e. in the absence of the applied field and dissipation, $\alpha=0;\,\,\eta=0$. Then for $\tau=0$
${\left.\left(\tilde{u}_{2}-u_{2}\right)\right|_{\tau=0}=\frac{1}{2}\left[\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right]}$
${=\frac{1}{2}\left[\,\frac{3}{2}u_{2}+\,\frac{1}{2}u_{2}\sqrt{1-\frac{8}{u_{2}^{2}\kappa^{2}}}-2u_{2}\right]=\frac{u_{2}}{4}\left[\,\sqrt{1-\frac{8}{u_{2}^{2}\kappa^{2}}}-1\right]}$ (3.15)
There arises an evident limitation from (III) (see [1]): $\frac{8}{u_{2}^{2}\kappa^{2}}<1;$ or $\varepsilon<\frac{u_{2}}{4}$. It is, however, not a severe limitation, it usually presumed $\varepsilon\ll 1$. Calculating $F$ given by (3.14), we get
@@
I.e. the correction due to memory has always the same sign as velocity (for positive $\kappa$ ); its relative importance increases with velocity and $\varepsilon$ (increasing $\varepsilon$ means decreasing steepness of the front).
Now let us consider the convective CHS equation: the field is non-zero, $\alpha\neq 0;$ but viscosity is still zero, $\eta=0$. It is worth noting that there is an upper limit on $\alpha$ in this case, corresponding to $s_{0}=0$:
@@
If $\alpha\to\alpha_{l}$, the $\tilde{u}_{1}$ root of (3.9) (unstable maximum of the potential) approaches zero (stable minimum); it means merging of stable and unstable states in the thermodynamic potential. So, the combined transition front is impossible for $\alpha>\alpha_{l}$.
@@
$F=\frac{\left(1-2\frac{u_{1}}{u_{2}}\right)-\alpha\kappa}{2\kappa^{2}\left[\sqrt{1+8\left(\frac{\alpha}{\kappa u_{2}^{2}}-\frac{1}{\kappa^{2}u_{2}^{2}}\right)}-1\right]}\left\{1+\frac{4\alpha\kappa-1}{\sqrt{1+8\left(\frac{\alpha}{\kappa u_{2}^{2}}-\frac{1}{\kappa^{2}u_{2}^{2}}\right)}}\right\}.$ (3.18)
Numeric example for (3.18) is given in Fig.1. The equation (3.18) could be simplified further. First, $\alpha\gg\frac{1}{\kappa}\sim\varepsilon\ll 1$. Then, according to (3.17) always $\alpha<\alpha_{l}$; so we can presume that $\alpha$ is not too close to its upper limit (which corresponds to extremely asymmetric potential), $\alpha\ll\alpha_{l}$, and $8\frac{\alpha}{\kappa u_{2}^{2}}=8\sqrt{2}\frac{\alpha\varepsilon}{u_{2}^{2}}\ll 1$. So, Eq. (3.18) is approximately
${F\simeq\frac{2}{\kappa^{2}}\left[\left(1-2\frac{u_{1}}{u_{2}}\right)-\alpha\kappa\right]\left(\frac{\kappa^{2}u_{2}^{2}+1}{4}-\alpha\kappa\right)}$
${=\frac{2}{\kappa^{2}}\left[\alpha\kappa-\left(1-2\frac{u_{1}}{u_{2}}\right)\right]\left(\alpha\kappa-\frac{\kappa^{2}u_{2}^{2}+1}{4}\right)}.$ (3.19)
Figure 1: Dependence of $F$ on parameters $\lambda=\alpha\kappa$ and $\rho=u_{1}/u_{2}$ according to (3.18). The values of $\kappa$ and $u_{2}$ are taken by way of example as $\kappa=10$ and $u_{2}=1$.
According to the latter expression the relative correction due to memory seems to be non-monotonic, has a minimum, and changes the sign at $\alpha\kappa=y_{1},y_{2}$:
$y_{1}=\left(1-2\frac{u_{1}}{u_{2}}\right);\,\,\,\,y_{2}=\frac{\kappa^{2}u_{2}^{2}+1}{4}.$ (3.20)
Per definition $-1<\left(1-2\frac{u_{1}}{u_{2}}\right)<1$; so $\alpha\kappa$ can cross $y_{1}$; it also means change of the velocity sign. However, according to our assumption, $\frac{\alpha}{\kappa u_{2}^{2}}\ll 1;$ i.e. $\alpha\kappa\ll\frac{1}{2}\left(y_{1}+y_{2}\right)\simeq\frac{1}{8}\kappa^{2}u_{2}^{2}$. That is, when $\alpha\kappa$ increases, the relative correction due to memory will be positive and having decreasing absolute value for $\alpha\kappa<y_{1}$, and negative with increasing absolute value for $y_{1}<\alpha\kappa$. I.e. the dependence is monotonic in the proper domain, see Fig.1, where $\alpha\kappa$ will not reach the position of minimum of (III).
Now let us consider the viscous CHS equation: the field is zero, $\alpha=0$ but the viscosity is non-zero, $\eta\neq 0$. Then from (3.4)-(3.5)
@@
If $\eta\to\eta_{l}$, the $\tilde{u}_{1}$ root of (3.9) approaches zero; it means again merging of stable and unstable states in the thermodynamic potential. So, the combined transition front is impossible for $\eta>\eta_{l}$.
Using (3.21) and (3.22) we get
$\delta_{0}^{2}+4s_{0}=\left[\frac{u_{2}}{2}+\left(\frac{u_{2}}{2}-u_{1}\right)\eta\right]^{2}-\frac{2}{\kappa^{2}}=\left(\frac{u_{2}}{2}+v\kappa\eta\right)^{2}-\frac{2}{\kappa^{2}}.$ (3.24)
Velocity (3.3) is not dependent on $\eta$, so the expressions (3.7) and (3.8) for $\delta_{1}$ and $s_{1}$ are the same, as in the case $\alpha=0;\,\,\eta=0$. Then
$\delta_{0}\delta_{1}+2s_{1}=-\left[\frac{1}{2}u_{2}+v\kappa\eta\right]\frac{v}{2\kappa},$ (3.25)
@@
I.e. for positive $v$ the effect of memory decreases monotonically with increasing viscosity.
IV Discussion
In the present work we studied the effect of memory on the model, introduced in [1]. It is a model of interplay of equilibrium and non-equilibrium phase transitions. Such interplay was analyzed by considering the advancing front which ‘‘combines’’, in some sense, these both transitions. While the equilibrium phase transition was described by the modified [2, 3] Cahn-Hilliard equation, the non-equilibrium phase transition was presented by the canonical chemical models introduced by Schlögl [4].
Schlögl considered two reaction systems: the so-called ‘‘First Schlögl Reaction’’ (1.8)-(1.9), which is an analog of the second order equilibrium phase transition, and the ‘‘Second Schlögl Reaction’’ (1.10)-(1.11), which is an analog of the first order equilibrium phase transition, for details see [4]. In these models the different ‘‘phases’’ correspond to different stationary states of the chemical reactions system.
To describe the equilibrium phase transition we used the Convective-Viscous Cahn-Hilliard equation; i.e. the classic Cahn-Hilliard equation complemented by the convective term due to the applied field, and by the viscous term to account for dissipation [11, 12, 13, 14, 2, 3].
The study of the classic Cahn-Hilliard equation complemented by the nonlinear sink/source terms was pioneered by Huberman [15] and Cohen and Murray [16]. It was followed by other papers of physical, biological and mathematical nature [17, 18, 19, 20, 21, 22]. Huberman [15] and Cohen and Murray [16], apparently unaware of Schlögl model, in fact considered interplay of equilibrium and second-order non-equilibrium phase transitions.
In [1] we have considered the convective-viscous Cahn-Hilliard equation with additional nonlinear terms, corresponding both to first and second Schlögl Reactions; there are quadratic and cubic nonlinearities, respectively. We obtained exact traveling waves solutions for these equations.
@@ IV Discussion
In [1] it was shown, that for the convective-viscous CH equation with first Schlögl reaction the conditions of simultaneous equilibrium and non-equilibrium transitions are very restrictive. Even more, equation for the traveling wave becomes degenerate in the absence of the applied field. On the other hand, for the CHS equation the influence of the non-equilibrium transition, i.e. of the reaction system, is much stronger. The transition front may be stopped, or even reversed both by changing the stationary states of the reaction system and by the field. The final state may be under-, or oversaturated, creating non-equilibrated phases. So, exploring the changes introduced by memory in the present paper we considered the convective-viscous CHS equation only.
In [1] description of the mass transfer was based on the ‘‘generalized Fick’s law’’ (1.3); in the present paper we use Maxwell-Cattaneo approach [23]. The exact traveling wave solution (2.18) has the same form, as in [1]. Even more, the expressions for the parameters of solution are exactly the same. However, the compatibility conditions (2.15), (2.17) are changed. The additional terms are proportional to non-dimensional characteristic time $\tau$. In dimensional (dimensionless) parameters the expression for $\tau$ is
@@
Here $\tau$ is the ratio of the initial characteristic time in Maxwell-Cattaneo equation to the characteristic time of the chemical reaction; it is usually considered to be a small parameter, $\tau\ll 1$. So, considering the expressions for the stationary states of thermodynamic potential (3.10), we can always expand them in $\tau$ and keep the linear terms only, see(3.11). We are interested in changes, due to memory, of the deviations of actual stationary concentration after transition from its equilibrium value. To measure this relative changes we introduced the function $F$, see (3.14). Of course, the relative change is $\tau F$, but we will drop $\tau$ herein.
Proceeding along the same lines as in [1], we considered three special cases. First we considered the CHS case, when the applied field and the dissipation are absent, $\alpha=0;\,\,\eta=0$, i.e. the classic CH equation complemented by the cubic nonlinearity. The single limitation in this case is $4\varepsilon<u_{2}$, or
@@
Here $\bar{\varepsilon}$ is of the order of capillarity length; $L$ is the diffusion distance for the characteristic time of the chemical reaction; so this limitation is not restrictive. From (3.16) it is evident, that the relative value of memory correction is zero for zero velocity of the front and is maximal for $u_{1}=0,$ that is when unstable state of chemical system merges with the stable state $u=0$.
If $\alpha\neq 0;\,\,\eta=0$, there is an upper limit on $\alpha$, see(3.17), neglecting small addition
@@
So, $\alpha_{l}$ is rather large; but if $\alpha\to\alpha_{l}$ , the $\tilde{u}_{1}$ root of (3.9) approaches zero; it means merging of stable and unstable states in the thermodynamic potential. So, the combined transition front is impossible for $\alpha>\alpha_{l}$. The approximate expression (III) for $F$, as well as numerical example in Fig.1, shows that with growing $\alpha\kappa$ the relative correction is positive and decreases for positive velocity, is zero for zero velocity and is negative with increasing absolute value for negative velocity. So, the memory effect became more pronounced, when the field forces the ‘‘retreat’’ of the front, acting against the normal direction determined by the chemical system.
Considering the viscous CHS equation we have $\eta\neq 0;\,\,\alpha=0$. Similar to the case $\alpha\neq 0$, there is an upper limit on $\eta$ in this case (3.23), corresponding to merging of unstable $\tilde{u}_{1}$ and stable $\tilde{u}=0$ states of the thermodynamic potential. Neglecting small correction, this limit is
$\eta_{l}\simeq\frac{u_{2}}{u_{2}-2u_{1}}$ (4.4)
for the velocity approaching zero $\eta_{l}\to\infty$; however the exact expression (3.26) gives zero correction. For positive velocity the correction decreases monotonically with viscosity. For negative velocity there is an additional limitation: if $\eta\to\eta_{m}$,
$\eta_{m}=\frac{\kappa u_{2}-2\sqrt{2}}{2\kappa u_{1}-\kappa u_{2}}$ (4.5)
the unstable state $\tilde{u}_{1}$ of the potential merges with stable state $\tilde{u}_{2}$, see (3.9) and the model becomes physically senseless. However, for $\eta<\eta_{m}$ the absolute value of the correction increases with viscosity $\eta$, and for $\eta>\eta_{m}$ decreases.
Summing up, for positive velocity of the transition front, when the state with larger concentration advances, both the applied field and viscosity generally diminish the memory corrections. For negative velocity the consistency of the model imposes stronger limitations on the allowed intervals of the parameters.
Acknowledgment
We are thankful to O.S. Bakai for the attention to the paper and valuable remarks.

### v2

@@ 3 The parametric dependence of solution
Naturally, for $\tau=0$, the equations (2.15), (2.17) coincide with the corresponding equations of Part I [1] too. We denote the values, corresponding to $\tau=0$ as $\delta_{0}$ and $s_{0}$:
$\delta_{0}=\,\,\frac{3}{2}u_{2}-\left[\left(\frac{u_{2}}{2}-u_{1}\right)-\kappa\frac{\alpha u_{2}}{2}\right]\eta,$ (3.4)
@@
$\delta=\delta_{0}+\tau\delta_{1},\quad s=s_{0}+\tau s_{1}.$ (3.6)
@@
Naturally, the corrections due to the memory are proportional to the velocity of the front; there should be no corrections for a static front.
The constraints (2.15) and (2.17) impose some limitations on the parameters $\delta$ and $s$ of the equilibrium phase transition, see (1.2), i.e., on the roots $\tilde{u}_{1}$, $\tilde{u}_{2}$, $\tilde{u}_{3}$ of the cubic equation
@@
These roots are positions of the extrema of the homogeneous part of the thermodynamic potential $\Phi\left(u\right)$; we denote the roots, corresponding to stable minima $\,\tilde{u}_{2}$, $\tilde{u}_{3}$, and $\tilde{u}_{1}$ — corresponding to an unstable maximum. The root $\tilde{u}_{3}=0$ coincides with one of the stationary states of the reaction system. The expressions for two remaining roots $\tilde{u}_{1}$ and $\tilde{u}_{2}$ yield the constraints imposed on the stationary values for the equilibrium transformation:
@@
We are mainly interested in what extent the actual final concentration $u_{2}$ after the transition deviates from the equilibrium value $\,\tilde{u}_{2}$. Usually $\tau$ is assumed to be a small parameter; expanding (3.10) and keeping only linear in $\tau$ terms, we have got
@@
Then, the deviation of the equilibrium value from the actual one after the transition is
$\tilde{u}_{2}-u_{2}=\frac{1}{2}\left[\left(\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right)+\left(\delta_{1}+\frac{\delta_{0}\delta_{1}+2s_{1}}{\sqrt{\delta_{0}^{2}+4s_{0}}}\right)\tau\right].$ (3.12)
The $\tau$-independent term was calculated and discussed in [1]; here, we are interested in the relative importance of the memory effect
${\tilde{u}_{2}-u_{2}=\frac{1}{2}\left(\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right)}$
×${\times\left[1+\left(\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right)^{-1}\left(\delta_{1}+\frac{\delta_{0}\delta_{1}+2s_{1}}{\sqrt{\delta_{0}^{2}+4s_{0}}}\right)\tau\right]}.$ (3.13)
Thus, we define
$F=\left(\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right)^{-1}\left(\delta_{1}+\frac{\delta_{0}\delta_{1}+2s_{1}}{\sqrt{\delta_{0}^{2}+4s_{0}}}\right).$ (3.14)
We study the behavior of $F$ as a function of the system parameters. Following [1], we consider several special cases. First we consider CHS, i.e., (2), where the applied field and dissipation are absent, $\alpha=0;\,\,\eta=0$. Then, for $\tau=0$ we have
${\left.\left(\tilde{u}_{2}-u_{2}\right)\right|_{\tau=0}=\frac{1}{2}\left(\delta_{0}+\sqrt{\delta_{0}^{2}+4s_{0}}-2u_{2}\right)}$
${=\frac{1}{2}\left(\frac{3}{2}u_{2}+\,\frac{1}{2}u_{2}\sqrt{1-\frac{8}{u_{2}^{2}\kappa^{2}}}-2u_{2}\right)=\frac{u_{2}}{4}\left(\sqrt{1-\frac{8}{u_{2}^{2}\kappa^{2}}}-1\right)}.$ (3.15)
From (3) an evident limitation arises (see [1]): ${8}/{u_{2}^{2}\kappa^{2}}<1;$ or $\varepsilon<{u_{2}}/{4}$. It is, however, not a severe limitation, it is usually assumed $\varepsilon\ll 1$. Calculating $F$ given by (3.14), we get
@@
I.e., the correction due to the memory has always the same sign as velocity (for positive $\kappa$); its relative importance increases with velocity and $\varepsilon$ (increasing $\varepsilon$ means decreasing steepness of the front).
Now, let us consider the convective CHS equation: the field is non-zero, $\alpha\neq 0;$ but the viscosity is still zero, $\eta=0$. It is worth noting that there is an upper limit on $\alpha$ in this case, corresponding to $s_{0}=0$:
@@
If $\alpha\to\alpha_{l}$, the $\tilde{u}_{1}$ root of (3.9) (unstable maximum of the potential) approaches zero (stable minimum); it means merging of stable and unstable states in the thermodynamic potential. Thus, the combined transition front is impossible for $\alpha>\alpha_{l}$.
@@
$F=\frac{\left(1-2\frac{u_{1}}{u_{2}}\right)-\alpha\kappa}{2\kappa^{2}\left[\sqrt{1+8\left(\frac{\alpha}{\kappa u_{2}^{2}}-\frac{1}{\kappa^{2}u_{2}^{2}}\right)}-1\right]}\left[1+\frac{4\alpha\kappa-1}{\sqrt{1+8\left(\frac{\alpha}{\kappa u_{2}^{2}}-\frac{1}{\kappa^{2}u_{2}^{2}}\right)}}\right].$ (3.18)
Numeric example for (3.18) is given in figure 1. The equation (3.18) could be simplified further. First, $\alpha\gg\frac{1}{\kappa}\sim\varepsilon\ll 1$. Then, according to (3.17) always $\alpha<\alpha_{l}$; so we can presume that $\alpha$ is not too close to its upper limit (which corresponds to extremely asymmetric potential), $\alpha\ll\alpha_{l}$, and $8({\alpha}/{\kappa u_{2}^{2}})=8\sqrt{2}({\alpha\varepsilon}/{u_{2}^{2}})\ll 1$. Thus, equation (3.18) is approximately
$F\simeq\frac{2}{\kappa^{2}}\left[\alpha\kappa-\left(1-2\frac{u_{1}}{u_{2}}\right)\right]\left(\alpha\kappa-\frac{\kappa^{2}u_{2}^{2}+1}{4}\right).$ (3.19)
Figure 1: (Colour online) Dependence of $F$ on parameters $\lambda=\alpha\kappa$ and $\rho=u_{1}/u_{2}$ according to (3.18). The values of $\kappa$ and $u_{2}$ are taken by way of example as $\kappa=10$ and $u_{2}=1$.
According to the latter expression, the relative correction due to the memory seems to be non-monotonous, has a minimum, and changes the sign at $\alpha\kappa=y_{1},y_{2}$:
$y_{1}=\left(1-2\frac{u_{1}}{u_{2}}\right);\quad y_{2}=\frac{\kappa^{2}u_{2}^{2}+1}{4}.$ (3.20)
Per definition $-1<1-2\frac{u_{1}}{u_{2}}<1$; so $\alpha\kappa$ can transverse $y_{1}$; it also means the change of the velocity sign. However, according to our assumption, ${\alpha}/{\kappa u_{2}^{2}}\ll 1;$ i.e. $\alpha\kappa\ll\frac{1}{2}\left(y_{1}+y_{2}\right)\simeq\frac{1}{8}\kappa^{2}u_{2}^{2}$. That is, when $\alpha\kappa$ increases, the relative correction due to the memory will be positive and will have a decreasing absolute value for $\alpha\kappa<y_{1}$, and will be negative with an increasing absolute value for $y_{1}<\alpha\kappa$. I.e., the dependence is monotonous in the proper domain, see figure 1, where $\alpha\kappa$ will not reach the position of minimum of (3.19).
Now, let us consider the viscous CHS equation: the field is zero, $\alpha=0$ but the viscosity is non-zero, $\eta\neq 0$. Then, from (3.4)–(3.5)
@@
If $\eta\to\eta_{l}$, the $\tilde{u}_{1}$ root of (3.9) approaches zero; it means again the merging of stable and unstable states in the thermodynamic potential. Thus, the combined transition front is impossible for $\eta>\eta_{l}$.
Using (3.21) and (3.22), we get
$\delta_{0}^{2}+4s_{0}=\left(\frac{u_{2}}{2}+v\kappa\eta\right)^{2}-\frac{2}{\kappa^{2}}.$ (3.24)
Velocity (3.3) is not dependent on $\eta$, so the expressions (3.7) and (3.8) for $\delta_{1}$ and $s_{1}$ are the same as in the case $\alpha=0$; $\eta=0$. Then,
$\delta_{0}\delta_{1}+2s_{1}=-\left(\frac{1}{2}u_{2}+v\kappa\eta\right)\frac{v}{2\kappa}.$ (3.25)
@@
I.e., for positive $v$ the effect of the memory decreases monotonously with an increasing viscosity.
4 Discussion
In the present work we studied the effect of the memory on the model, introduced in [1]. It is a model of interplay of equilibrium and non-equilibrium phase transitions. Such an interplay was analyzed by considering the advancing front which “combines”, in some sense, these both transitions. While the equilibrium phase transition was described by the modified [2, 3] Cahn–Hilliard equation, the non-equilibrium phase transition was presented by the canonical chemical models introduced by Schlögl [4].
In [1] we considered the convective-viscous Cahn–Hilliard equation with additional nonlinear terms, corresponding both to the first and second Schlögl reactions; there are quadratic and cubic nonlinearities, respectively. We obtained the exact travelling waves solutions for these equations.
@@ 4 Discussion
As in [1], to make the model formulation possibly transparent, we assumed the last reaction of the system to be irreversible; indeed, by changing the kinetic parameter and/or diminishing the concentration $C$, this reaction could always be shifted close to irreversible. On the other hand, for the reversible reaction, the equilibrium concentration before transition will also deviate from the actual one; this will demand a special preparation of the system and make the model quite artificial.
In [1] the description of the mass transfer was based on the “generalized Fick’s law” (1.3); in the present paper we introduce the memory effect, i.e., we use Maxwell-Cattaneo approach (1.9) [23]. The most important result of the present work is that despite the introduction of the memory effects, the CHS equation still possesses an exact travelling wave solution. This is nontrivial: many well known diffusive model equations, e.g., Newell-Whitehead equation, Nagumo equation, Zeldovich equation etc., see [30], lose this property, when the memory effects are introduced, see [31].
The exact travelling wave solution is a “tool”, which exhibits clearly the effect of the memory. The exact solution (2.18) has the same form, as in [1]. Even more, the expressions for the parameters of solution are exactly the same. I.e., while for the non-stationary processes the memory effect can change the rates and the form of the wave, see, e.g., [23], for the constant velocity transition, the velocity and steepness of the wave are the same. However, the compatibility conditions (2.15), (2.17) are changed. This changes correspondingly the possible over/undersaturation after the transition. The additional terms are proportional to the non-dimensional characteristic time $\tau$. In dimensional parameters, the expression for $\tau$ is
@@
Here, $\tau$ is the ratio of the initial characteristic time in Maxwell-Cattaneo equation to the characteristic time of the chemical reaction; it is usually considered to be a small parameter, $\tau\ll 1$. Thus, considering the expressions for the stationary states of thermodynamic potential (3.10), we can always expand them in $\tau$ and keep the linear terms only, see (3.11). We are interested in the changes, due to memory, of the deviations of the actual stationary concentration after transition from its equilibrium value. To measure these relative changes, we introduced the function $F$, see (3.14). Of course, the relative change is $\tau F$, but we will drop $\tau$ herein.
Proceeding along the same lines as in [1], we considered three special cases. First we considered the CHS case when the applied field and the dissipation are absent, $\alpha=0;\,\,\eta=0$, i.e., the classic CH equation complemented by the cubic nonlinearity. The single limitation in this case is $4\varepsilon<u_{2}$, or
@@
Here, $\bar{\varepsilon}$ is of the order of capillarity length; $L$ is the diffusion distance for the characteristic time of the chemical reaction; so this limitation is not restrictive. From (3.16) it is evident, that the relative value of memory correction is zero for zero velocity of the front and is maximal for $u_{1}=0,$ that is when unstable state of the chemical system merges with the stable state $u=0$. It is also increasing with $\varepsilon$; increasing $\varepsilon$ means decreasing steepness of the front. Indeed, for the fast moving shallow front, the relaxation to equilibrium is delayed, and the integrated over $\tau$ influence of the “history” increases.
If $\alpha\neq 0$; $\eta=0$, there is an upper limit on $\alpha$, see (3.17), neglecting a small addition
@@
Thus, $\alpha_{l}$ is rather large; but if $\alpha\to\alpha_{l}$, the $\tilde{u}_{1}$ root of (3.9) approaches zero; it means merging of stable and unstable states in the thermodynamic potential. Thus, the combined transition front is impossible for $\alpha>\alpha_{l}$. The approximate expression (3.19) for $F$, as well as numerical example in figure 1, show that with growing $\alpha\kappa$ the relative correction is positive and decreases for positive velocity, it is zero for zero velocity and it is negative with increasing absolute value for negative velocity. Thus, the memory effect became more pronounced, when the field forces the “retreat” of the front, acting against the normal direction determined by the chemical system. This again leads to a delayed relaxation.
Considering the viscous CHS equation, we have $\eta\neq 0$; $\alpha=0$. Similar to the case $\alpha\neq 0$ , there is an upper limit on $\eta$ in this case, (3.23), corresponding to the merging of unstable $\tilde{u}_{1}$ and stable $\tilde{u}=0$ states of the thermodynamic potential. Neglecting small correction, this limit is as follows:
$\eta_{l}\simeq\frac{u_{2}}{u_{2}-2u_{1}}.$ (4.4)
For the velocity approaching zero $\eta_{l}\to\infty$, although the exact expression (3.26) gives zero correction. For positive velocity, the correction decreases monotonously with increasing viscosity. For negative velocity, there is an additional limitation: if $\eta\to\eta_{m}$,
$\eta_{m}=\frac{\kappa u_{2}-2\sqrt{2}}{2\kappa u_{1}-\kappa u_{2}},$ (4.5)
the unstable state $\tilde{u}_{1}$ of the potential merges with stable state $\tilde{u}_{2}$, see (3.9), and the model becomes physically senseless. However, for $\eta<\eta_{m}$, the absolute value of the correction increases with viscosity $\eta$, and for $\eta>\eta_{m}$, it decreases.
Summing up, for positive velocity of the transition front, when the state with larger concentration advances, both the applied field and viscosity generally diminish the memory corrections. For negative velocity, the consistency of the model imposes stronger limitations on the allowed intervals of the parameters; the delayed relaxation makes the memory effects more pronounced. I.e., if the diffusion in the given system demonstrates the memory effects, changing the transition velocity, and the applied field yields an additional instrument to achieve the necessary over/undersaturation after transition.
Acknowledgements
We are thankful to O. S. Bakai for the attention to the paper and for valuable remarks.